A worker is exposed to:

90 dBA of noise for 4 hrs and 45 minutes,

92 dBA for 2 hrs and 15 minutes, and 9

5 dBA of noise for 1 hr.

a). What is the value of TWAN ? __ __

b). Is the worker overexposed to noise (Yes or No)?
Answer #1

a)

The formula used for Dose

D = 100*(C1/T1 + C2/T2 +……Cn/Tn)

Here, we need to calculate the T values from the Tn Lookup table. We can get this from OSHA regulation documents/website. We can also use the formula T = 8/2^((L-90)/5) where L is the sound level. The values of C is the number of hours. So we have

C1 = 4.75, C2 = 2.25, C3 = 1

T1 = 8, T2 = 6.1, T3 = 4

D = 100*(4.75/8 + 2.25/6.1 + 1/4) = 121.26

Again we can use the TWA lookup table. However, we can also calculate it using the formula

TWA = 16.61 Log10(D/100) + 90

Using this we can calculate

TWA = 16.61*Log10(121.26/100)+90 = 91.39

b)

The OSHA standards allow for TWA <= 90.

Considering that the worker is exposed to 91.39, he/she is overexposed